OpenStudy (anonymous):

solve the system of equations: w/4+z/6=4 w/2-z/3=4

OpenStudy (texaschic101):

w/4 + z/6 = 4 (multiply by 12) 3w + 2z = 48 w/2 - z/3 = 4 (multiply by 6) 3w - 2z = 24 now we have : 3w + 2z = 48 3w - 2z = 24 ------------ 6w = 72 w = 12 3w + 2z = 48 3(12) + 2z = 48 36 + 2z = 48 2z = 48 - 36 2z = 12 z = 6 check... w/4 + z/6 = 4 12/4 + 6/6 = 4 36/12 + 12/12 = 4 48/12 = 4 4 = 4 (correct) w = 12 and z = 6